My son is in 10th grade, taking Algebra II. They just started this new section last week when my son was out sick two days. He is struggling with understanding his make-up work and won't be able to see the teacher until Tuesday after school. I hope you can help me help him in the meantime?
I'm not sure how to symbolize square root in your system either?
I'll use "sq rt" for square root -- ok?
Here is what he's done so so far, squaring, simplifying and getting the GCF
(5 times sq rt of 3 + 3 times sq rt of 5)(squared)
work:
(25x9) + 30 times sq rt of 15 + (9x5)
answer:
270 + 30 times sq rt of 15
Is that anywhere near correct?
Thank you so much!

Answer by stanbon(48570)

(5 times sq rt of 3 + 3 times sq rt of 5)(squared)
work:
(25x9) + 30 times sq rt of 15 + (9x5)
answer:
270 + 30 times sq rt of 15
-------------
(5sqrt3 + 3sqrt5)
= (5sqrt3)^2 +2(5sqrt3)(3sqrt5) + (3sqrt5)^2
= 25*3 + 2*15*sqrt15 + 9*5
= 75 +30sqrt15 + 45
= 120 + 30sqrt15
----------------------
Cheers,
stan H.